# Global decomposition of networks into multiple cores formed by local hubs

### v2

@@ II.3 Local-edge decomposition of the Barabási-Albert model
Figure 4: (a) The giant component size $G$ as a function of the fraction $p$ of removed edges based on $h\mbox{-}\mathrm{P}$, applied to a single realization of the BA model [44] with $10^{5}$ nodes and the number $m=4$ of stubs for each newly added node. The black squares indicate the original network. The red circles, blue triangles, and purple inverted triangles represent the results obtained for the primary, secondary, and tertiary backbones, respectively, of the original network. (b) The relative fraction $g$ of the giant component size with zero hub-centrality nodes removed as a function of the fraction $n_{0}$ of zero hub-centrality nodes in decomposed networks at different levels. The black squares, red circles, and blue triangles represent the cases with $m=2$, $4$, and $6$, respectively. We only plot the cases where the backbone with the number of nodes $>1\%$ of that of the original network.
@@ II.4 Comparison with the k-core decomposition
Figure 5: The fraction of the level (or “shell” in the kCD) relationship between nodes and their neighbors. The colors red, blue, and green indicate that nodes at a given level have neighbors with higher, lower, or the same level, respectively. The regions filled with gray represent the absence of nodes in the corresponding levels. The panel (a) is for the LED, and the panel (b) is for the kCD.
@@
Figure 6: A schematic illustration of examples of global and local hubs, whose horizontal location is aligned with their hub-centrality values represented with the horizontal color bar with gradation.
@@ III.1 Core-periphery structure and score
Figure 7: The connection densities of core-core, core-periphery, and periphery-periphery, along with the core-periphery score, as a function of the boundary level $L_{B}$. The black (square), red (circle), and green (triangle) points represent the connection densities of core-core, core-periphery, and periphery-periphery, respectively. The purple (inverse triangle) points indicate the core-periphery score. At $L_{B}=0$, the connection densities for core-periphery and periphery-periphery, as well as the core-periphery score, are zero. The core-periphery score reaches its maximum at $L_{B}=18$.
First, in order to demonstrate that our LED identifies the single CP, we observe the connection density of core-core ($D_{cc}=2M_{cc}/[N_{c}(N_{c}-1)]$), core-periphery ($D_{cp}=M_{cp}/(N_{c}N_{p})$), and periphery-periphery ($D_{pp}=2M_{pp}/[N_{p}(N_{p}-1)]$) edges, where $N_{c}$, $N_{p}$, $M_{cc}$, $M_{cp}$, and $M_{pp}$ correspond to the number of core nodes, that of periphery nodes, that of edges connecting core nodes, that of edges connecting core and periphery nodes, and that of edges connecting periphery nodes, respectively. The conventional notion of CP suggests the inequality [14, 15, 16, 17, 18]
@@
    All of the periphery nodes are connected to core nodes, and no edges between periphery nodes exist.
@@
Basically, it represents an adjacency matrix with the perfect shape. The condition $1$ is quantified with $D_{cc}$, and the condition $2$ is measured by $M_{cp}/N_{p}$. The quantity $M_{cp}$, the number of edges connecting the core and periphery nodes, is also the number of periphery nodes connected to core nodes for a simple network without the possibility of multiple edges between a pair of nodes, so the ratio $M_{cp}/N_{p}$ represents the fraction of periphery nodes that are connected to the core nodes. In other words, the extent to which the core and periphery nodes satisfy the above two conditions is determined by $D_{cc}$ and $M_{cp}/N_{p}$, respectively. To quantitatively assess the core-periphery structure using these criteria, we define the core-periphery score as,
$S_{cp}\equiv D_{cc}\frac{M_{cp}}{N_{p}}-D\,\mathrm{min}\{1,DN_{c}\}\,,$ (3)
@@
where $D=2(M_{cc}+M_{cp}+M_{pp})/[(N_{c}+N_{p})(N_{c}+N_{p}-1)]$ is the density of the entire network. This represents the difference between the actual similarity and the expected similarity by chance as the null-model case, concerning the ideal CP. The core-periphery score is maximized when both conditions are satisfied, thus it helps to identify the most optimal boundary level $L_{B}^{*}$ for distinguishing between the core and periphery of a network. When a network exhibits an ideal CP structure, $S_{cp}$ is close to $1$. In contrast, in a null model network, $S_{cp}=0$ because the core-periphery score represents the comparison of the network with the null model by definition. As the boundary level $L_{B}$ increases from $1$ to $18$ ($S_{cp}$ for $L_{B}=0$ is undefined), the distinction between core and periphery nodes becomes clearer, causing the core-periphery score to gradually increase. At $L_{B}^{*}=18$, the core-periphery score reaches its maximum value, and this means that nodes belonging to $L\geq L_{B}^{*}=18$ form the core, while the remaining nodes form the periphery in this network. However, when $L_{B}$ exceeds $L_{B}^{*}=18$, $M_{cp}$ decreases and the core-periphery score starts to decrease.
Despite the existence of optimal boundary $L_{B}^{*}=\mathrm{arg\,max}_{L_{B}}S_{cp}$, the core-periphery score itself from this dichotomous distinction between a single core and the rest as the periphery is quite small ($\lesssim 0.1$) even near $L_{B}^{*}$ in the CB network, as shown in Fig. 7. As previously discussed, it stems from the obvious fact that the assumption of single core-periphery separation has its clear limitation for this large-scale real network. The macroscale network is composed of various mesoscale structural features (multiple hubs, communities, etc.) and their interwoven mixture. Since the LED uses nodes’ local relative-degree information, it simultaneously decomposes the network in various heterogeneous places, e.g., communities with different scales, as discussed in Sec. II.4 with Fig. 6, which causes the low value of $S_{cp}$; both $D_{cc}$ and $N_{cp}/N_{p}$ cannot have high values simultaneously.
@@ III.2 Core-periphery structure of communities and the supernode network
Figure 8: (a) An example of a network with a community structure marked by different background colors. (b) Inside an arbitrary community of the CB network, we plot the connection densities of core-core, core-periphery, and periphery-periphery, along with the core-periphery score, as a function of the boundary level $L_{B}$. The same colors and shapes are used for the densities and the score as in Fig. 7. The core-periphery score $S_{cp}$ reaches its maximum at $L_{B}=9$. (c) The adjacency matrix represents the connections (yellow in the black background) inside the community in panel (b), where nodes are sorted by their LED levels (higher levels toward the left bottom). The nodes with the lowest level are excluded for visibility, and the nodes inside the red square are the optimal set of core nodes of this community.
Figure 9: (a) The supernode network is composed of the nodes corresponding to the communities in the network shown in Fig. 8(a). (b) The connection densities of core-core, core-periphery, and periphery-periphery, along with the core-periphery score, as a function of the boundary level $L_{B}$ in the supernode network from the CB network. The same colors and shapes are used for the densities and the score as in Figs. 7 and 8. The core-periphery score $S_{cp}$ reaches its maximum at $L_{B}=4$. (c) The adjacency matrix represents the connections (green in the black background) between the supernodes in panel (b), where nodes are sorted by their LED levels (higher levels toward the left bottom). The supernodes inside the red square are the optimal core nodes of this community.
In the previous subsection, we have found a single core of the network, but as introduced in the first part of Sec. III, more recent studies on the CP emphasize the necessity for the consideration of multiple CP in networks [19, 20, 21, 22]. One hint from the literature is the fact that the mixture of community structures and CP is interchangeably expressed as ‘CP inside communities’ and ‘communities inside CP’ (see Fig. 1.1 of Ref. [17]). In this subsection, we take both viewpoints by applying the LED to the nodes inside each community (the former) and to coarse-grained communities (the latter). For this (literally) divide-and-conquer strategy, we use the Louvain method [53, 54] to systematically obtain community structures. With our local-hub-based LED scheme applied to each community, we expect to identify core backbone nodes and peripheral shell nodes, which would be closely related to individual nodes’ co-membership consistency with other nodes from stochastic community-detection algorithms (consistent core nodes versus inconsistent peripheral nodes) [55, 56].
@@
To check the ‘communities inside CP’ side, we take the coarse-grained point of view and treat each community detected by the Louvain algorithm [53, 54] as a new node or a “supernode” for distinction. The supernode network is composed of supernodes and the edges connecting them if there is any connection between the members of each community, as illustrated in Fig. 9(a) compared to Fig. 8(a); multi-edges and self-loops are ignored for simplicity. Then, on this coarse-grained network [59, 60], we apply the same LED process to detect the CP. Figure 9(b) shows the result for the supernode network from the CB network, again in the case of $\gamma=0.5$. The result indicates $L_{B}^{*}=4$ in this case, which assigns the $7$-clique in the super-adjacency matrix depicted in Fig. 9(c). The fact that $L_{B}^{*}$ corresponds to the final stage of decomposition may imply the lack of a characteristic scale or the lack of detailed resolution for this small-sized supernode network composed of $36$ supernodes; the latter hypothesis is supported by the same phenomenon happening for the LED inside a smaller community than the one depicted in Fig. 8. The super-adjacency matrix looks similar to the ideal CP form indeed; the core nodes are fully connected and most periphery nodes are connected to the core nodes. As in the inside-community version of LED, the structural simplification performed by coarse-graining with Louvain communities enables us to find a simple CP. There are communities or supernodes playing a core role in the super-network composed of supernodes. When using the Louvain method, the number of communities varies depending on the resolution parameter $\gamma$, and the nodes comprising each community change for each run even for a single $\gamma$ value due to its stochasticity. However, except for extreme cases where the resolution is very close to 00 or significantly greater than $1$, results are qualitatively similar to the result obtained in this section.
@@
In addition, we would like to remark on the utility of the hub centrality in this context. The degree as the number of neighbors measures a node’s global significance within the entire network, which is certainly a useful and the most widely used one. However, depending on the purpose of the analysis, comparing the (global) degree values across different focused groups can make deciphering hidden organizational principles even harder, especially in the case of quite heterogeneous degree distributions characterizing real networks [1, 2, 3]. The hub centrality [23, 7] is a tailor-made measure to quantify each node’s relative importance within each focused group composed of the node itself and its neighbors, and it can help us overcome such limitations by standardizing the status of local hubs (recall Fig. 6). Utilizing the hub centrality, the LED assigns standardized levels to local units and identifies the CP based on these levels, and thus, it is versatile across different scales, such as a network, within a community, or a supernode network.
@@ III.3 LED versus kCD for a synthetic network with a prescribed CP structure
Figure 10: (a) The number of core nodes and that of all nodes in the remained network with $L\geq L_{B}$ as a function of the boundary level $L_{B}$, from a single realization of the SBM structure in Sec. III.3. The squares and circles represent the results for the LED and kCD, respectively. The filled and open symbols indicate the core and all nodes, respectively. (b) The core-periphery score as a function of the boundary level. The filled squares and circles represent the cases of LED and kCD, respectively.
@@
In this final subsection, we demonstrate the LED’s merit even for a simple structure with gradual coreness. To show a clear advantage of the LED over kCD in this case, we generate a synthetic network using the stochastic block model (SBM) [61]. The synthetic network is composed of three parts denoted by $I_{1}$, $I_{2}$, and $I_{3}$, which are composed of $20$, $40$, and $100$ nodes, respectively. The connection density between different parts is given by $D_{I_{1}I_{1}}=0.7$, $D_{I_{2}I_{2}}=0.3$, $D_{I_{3}I_{3}}=0.1$, $D_{I_{1}I_{2}}=0.5$, $D_{I_{2}I_{3}}=0.2$, and $D_{I_{1}I_{3}}=0$; the nodes inside each group and between groups are connected uniformly at random with the prescribed connection density. The structure represents the three-level coreness, where the primary core part $I_{1}$ do not have any connection to the periphery $I_{3}$ and the secondary core $I_{2}$ plays the role of bridge between the primary core and the periphery. As shown in Fig. 10(a), in the case of kCD-based decomposition, the number of nodes with $L\geq L_{B}$ in the network does not notably change when $L_{B}<10$. In other words, even the level of most peripheral nodes typically goes up to higher than $L_{B}=10$, and only the $60$ nodes constituting the primary and secondary cores remain when $L_{B}\geq 16$. It means that the kCD roughly cuts the structure into two levels: both the primary and the secondary cores as the core and the rest as the periphery. When we apply the LED, in sharp contrast, the number of total nodes with $L\geq L_{B}$ decreases gradually, and in particular, the number of primary and secondary core nodes ($60$ nodes) gradually decreases for $L_{B}>10$, while most decrement in the number of nodes for $L_{B}<10$ is caused by the removal of peripheral nodes ($100$ nodes). Even if we set up the three-level coreness, due to the statistical fluctuation there must be differential levels of local hubs for a single realization of SBM, and the LED successfully captures them as well.
@@
We also measure $S_{cp}$ as a function of $L_{B}$, in the same manner as in Fig. 7. As shown in Fig. 10(b), in the case of the kCD, since most periphery nodes are in high levels ($>10$), $S_{cp}$ remains low across a wide range of lower $L_{B}$ values. At the maximum value of $L_{B}=19$, $S_{cp}$ reaches $0.320$ and the most densely connected $55$ nodes are classified as the core. However, in the case of the LED, $S_{cp}$ starts to increase and reaches its maximum value ($0.453$) at $L_{B}=L_{B}^{*}=16$, and decreases from that point. The number of nodes at $L_{B}^{*}$ is $20$ [Fig. 10(a)], which corresponds to the primary core nodes as expected. In other words, when using the LED, it is possible to identify the most influential nodes even within densely connected nodes. This behavior is consistent with that observed in real networks, shown in Figs. 7 and 8. As $L_{B}$ increases, the proportion of nodes belonging to $I_{1}$ increases among the nodes with $L>L_{B}$, leading to a reduction in the number of connections between nodes with $L\geq L_{B}$ and those with $L<L_{B}$. This results in a lower value of $N_{cp}/N_{p}$ and the presence of the optimal boundary level. Consequently, compared with the kCD, the LED allows the identification of more important nodes and a more detailed analysis even within densely connected structures, such as communities.

### v3

@@ II.3 Local-edge decomposition of the Barabási-Albert model
Figure 4: (a) Relative size $G$ of the giant component as a function of the fraction $p$ of removed edges based on $h\mbox{-}\mathrm{P}$, applied to a single realization of the BA model [11] with $10^{5}$ nodes and the number $m=4$ of stubs for each newly added node. The black squares indicate the original network. The red circles, blue triangles, and purple inverted triangles represent the results obtained for the primary, secondary, and tertiary backbones, respectively, of the original network. (b) Relative fraction $g$ of the giant component size with zero hub-centrality nodes removed as a function of the fraction $n_{0}$ of zero hub-centrality nodes in decomposed networks at different levels. The black squares, red circles, and blue triangles represent the cases with $m=2$, $4$, and $6$, respectively. We only plot the cases where the backbone with the number of nodes $>1\%$ of that of the original network.
@@ II.4 Comparison with the k-core decomposition
Figure 5: Fraction of the level (or “shell” in the kCD) relationship between nodes and their neighbors. The colors red, blue, and green indicate that nodes at a given level have neighbors with higher, lower, or the same level, respectively. The regions filled with gray represent the absence of nodes in the corresponding levels. Panel (a) is for the LED, and panel (b) is for the kCD.
@@
Figure 6: Schematic illustration of examples of global and local hubs, whose horizontal location is aligned with their hub-centrality values represented with the horizontal color bar with gradation.
@@ III.1 Core-periphery structure and score
Figure 7: Connection densities of core-core, core-periphery, and periphery-periphery, along with the core-periphery score, as a function of the boundary level $L_{B}$. The black (square), red (circle), and green (triangle) points represent the connection densities of core-core, core-periphery, and periphery-periphery, respectively. The purple (inverse triangle) points indicate the core-periphery score. At $L_{B}=0$, the connection densities for core-periphery and periphery-periphery, as well as the core-periphery score, are zero. The core-periphery score reaches its maximum at $L_{B}=18$.
First, to demonstrate that our LED identifies the single CP, we observe the connection density of core-core ($D_{cc}=2M_{cc}/[N_{c}(N_{c}-1)]$), core-periphery ($D_{cp}=M_{cp}/(N_{c}N_{p})$), and periphery-periphery ($D_{pp}=2M_{pp}/[N_{p}(N_{p}-1)]$) edges, where $N_{c}$, $N_{p}$, $M_{cc}$, $M_{cp}$, and $M_{pp}$ correspond to the number of core nodes, that of periphery nodes, that of edges connecting core nodes, that of edges connecting core and periphery nodes, and that of edges connecting periphery nodes, respectively. The conventional notion of CP suggests the inequality [14, 15, 16, 17, 18]
@@
    All of the periphery nodes are connected to at least a core node, and no edges between periphery nodes exist.
@@
The condition $1$ is quantified with $D_{cc}$, and the condition $2$ is measured by $N_{pc}/N_{p}$, where $N_{pc}$ is the number of periphery nodes with at least one connection to core nodes, so that the ratio $N_{pc}/N_{p}$ represents the fraction of periphery nodes that are connected to the core nodes. In other words, the extent to which the core and periphery nodes satisfy the above two conditions is determined by $D_{cc}$ and $N_{pc}/N_{p}$, respectively. To quantitatively assess the core-periphery structure using these criteria, we define the core-periphery score as,
$S_{cp}\equiv D_{cc}\frac{N_{pc}}{N_{p}}-D\,\mathrm{min}\{1,DN_{c}\}\,,$ (4)
@@
where $D=2(M_{cc}+M_{cp}+M_{pp})/[(N_{c}+N_{p})(N_{c}+N_{p}-1)]$ is the density of the entire network. This represents the difference between the actual similarity and the null-model case, where the same overall connection density $D$ applies for all potential connections. The second term on the right-hand side of Eq. (4) is the expected value of the first term for the case of completely random connections (the uniform edge density $D$ throughout the entire network) without any prescribed division between core and periphery nodes, as $D_{cc}=D$ and $N_{pc}/N_{p}=\mathrm{min}\{1,DN_{c}\}$ in that case; if $D>1/N_{c}$, for each periphery node there would be at least one edge to a core node on average, and if $D<1/N_{c}$, the expected fraction of periphery node with at least one connection to a core node would be $DN_{c}$. The core-periphery score is maximized when both conditions are satisfied, thus it helps to identify the most optimal boundary level $L_{B}^{*}$ for distinguishing between the core and periphery of a network. When a network exhibits an ideal CP structure, $S_{cp}$ is close to $1$. In contrast, in a null model network, $S_{cp}=0$ because the core-periphery score represents the comparison of the network with the null model by definition. As the boundary level $L_{B}$ increases from $1$ to $18$ ($S_{cp}$ for $L_{B}=0$ is undefined), the distinction between core and periphery nodes becomes clearer, causing the core-periphery score to gradually increase. At $L_{B}^{*}=18$, the core-periphery score reaches its maximum value, and this means that nodes belonging to $L\geq L_{B}^{*}=18$ form the core, while the remaining nodes form the periphery in this network. However, when $L_{B}$ exceeds $L_{B}^{*}=18$, $N_{pc}$ decreases and the core-periphery score starts to decrease.
In large-scale real networks, the assumption of single core-periphery separation has its clear limitation. Given that the macroscale network consists of various mesoscale structural features (multiple hubs, communities, etc.) and their interwoven mixture, the presence of multiple cores and a mesoscale core is an inherent characteristic of its structure.
@@ III.2 Core-periphery structure of communities and the supernode network
Figure 8: (a) Example of a network with a community structure marked by different background colors. (b) Inside an arbitrary community of the CB network, we plot the connection densities of core-core, core-periphery, and periphery-periphery, along with the core-periphery score, as a function of the boundary level $L_{B}$. The same colors and shapes are used for the densities and the score as in Fig. 7. The core-periphery score $S_{cp}$ reaches its maximum at $L_{B}=9$. (c) The adjacency matrix represents the connections (yellow in the black background) inside the community in panel (b), where nodes are sorted by their LED levels (higher levels toward the left bottom). The nodes with the lowest level are excluded for visibility, and the nodes inside the red square are the optimal set of core nodes of this community.
Figure 9: (a) Supernode network, which is composed of the nodes corresponding to the communities in the network shown in Fig. 8(a). (b) Connection densities of core-core, core-periphery, and periphery-periphery, along with the core-periphery score, as a function of the boundary level $L_{B}$ in the supernode network from the CB network. The same colors and shapes are used for the densities and the score as in Figs. 7 and 8. The core-periphery score $S_{cp}$ reaches its maximum at $L_{B}=4$. (c) The adjacency matrix represents the connections (green in the black background) between the supernodes in panel (b), where nodes are sorted by their LED levels (higher levels toward the left bottom). The supernodes inside the red square are the optimal core nodes of this community.
In the previous subsection, we have found a single core of the network, but as introduced in the first part of Sec. III, more recent studies on the CP emphasize the necessity for the consideration of multiple CP in networks [19, 20, 21, 22]. One hint from the literature is the fact that the mixture of community structures and CP is interchangeably expressed as “CP inside communities” and “communities inside CP” (see Fig. 1.1 of Ref. [17]). In this subsection, we take both viewpoints by applying the LED to the nodes inside each community (the former) and to coarse-grained communities (the latter). For this (literally) divide-and-conquer strategy, we use the Louvain method [53, 54] to systematically obtain community structures. With our local-hub-based LED scheme applied to each community, we expect to identify core backbone nodes and peripheral shell nodes, which would be closely related to individual nodes’ co-membership consistency with other nodes from stochastic community-detection algorithms (consistent core nodes versus inconsistent peripheral nodes) [55, 56].
@@
To check the “communities inside CP” side, we take the coarse-grained point of view and treat each community detected by the Louvain algorithm [53, 54] as a new node or a “supernode” for distinction. The supernode network is composed of supernodes and the edges connecting them if there is any connection between the members of each community, as illustrated in Fig. 9(a) compared to Fig. 8(a); multi-edges and self-loops are ignored for simplicity. Then, on this coarse-grained network [59, 60], we apply the same LED process to detect the CP. Figure 9(b) shows the result for the supernode network from the CB network, again in the case of $\gamma=0.5$. The result indicates $L_{B}^{*}=4$ in this case, which assigns the $7$-clique in the super-adjacency matrix depicted in Fig. 9(c). The fact that $L_{B}^{*}$ corresponds to the final stage of decomposition may imply the lack of a characteristic scale or the lack of detailed resolution for this small-sized supernode network composed of $36$ supernodes; the latter hypothesis is supported by the same phenomenon happening for the LED inside a smaller community than the one depicted in Fig. 8. The super-adjacency matrix looks similar to the ideal CP form that we discussed; the core nodes are densely connected and most periphery nodes are connected to core nodes. As in the inside-community version of LED, the structural simplification performed by coarse-graining with Louvain communities enables us to find a simple CP. There are communities or supernodes playing a core role in the super-network composed of supernodes. When using the Louvain method, the number of communities varies depending on the resolution parameter $\gamma$, and the nodes comprising each community change for each run even for a single $\gamma$ value due to its stochasticity. However, except for extreme cases where the resolution is very close to 00 or significantly greater than $1$, results are qualitatively similar to the result obtained in this section.
@@
In addition, we remark on the utility of the hub centrality in this context. The degree as the number of neighbors measures a node’s global significance within the entire network, which is certainly a useful and the most widely used one. However, depending on the purpose of the analysis, comparing the (global) degree values across different focused groups can make deciphering hidden organizational principles even harder, especially in the case of quite heterogeneous degree distributions characterizing real networks [1, 2, 3]. The hub centrality [23, 7] is a tailor-made measure to quantify each node’s relative importance within each focused group composed of the node itself and its neighbors, and it can help us overcome such limitations by standardizing the status of local hubs (recall Fig. 6). Utilizing the hub centrality, the LED assigns standardized levels to local units and identifies the CP based on these levels, and thus, it is versatile across different scales, such as a network, within a community, or a supernode network.
@@ III.3 LED versus kCD for a synthetic network with a prescribed CP structure
Figure 10: Connection density scheme of the stochastic block model, which is composed of three parts: $I_{1}$, $I_{2}$, and $I_{3}$. The relative length of the side is proportional to the number of nodes in each block.
@@
Figure 11: (a) Number of core nodes and that of all nodes in the remained network with $L\geq L_{B}$ as a function of the boundary level $L_{B}$, from a single realization of the SBM structure in Sec. III.3. The squares and circles represent the results for the LED and kCD, respectively. The filled and open symbols indicate the core and all nodes, respectively. (b) Core-periphery score as a function of the boundary level. The filled squares and circles represent the cases of LED and kCD, respectively.
@@
In this final subsection, we demonstrate the LED’s merit even for a simple structure with gradual coreness. To show a clear advantage of the LED over kCD in this case, we generate a synthetic network using the stochastic block model (SBM) [61]. The synthetic network is composed of three parts denoted by $I_{1}$, $I_{2}$, and $I_{3}$, which are composed of $20$, $40$, and $100$ nodes, respectively. As shown in Fig. 10, the connection density between different parts is given by $D_{I_{1}I_{1}}=0.7$, $D_{I_{2}I_{2}}=0.3$, $D_{I_{3}I_{3}}=0.1$, $D_{I_{1}I_{2}}=0.5$, $D_{I_{2}I_{3}}=0.2$, and $D_{I_{1}I_{3}}=0$; the nodes inside each group and between groups are connected uniformly at random with the prescribed connection density. The structure represents the three-level coreness, where the primary core part $I_{1}$ do not have any connection to the periphery $I_{3}$ and the secondary core $I_{2}$ plays the role of bridge between the primary core and the periphery. As shown in Fig. 11(a), in the case of kCD-based decomposition, the number of nodes with $L\geq L_{B}$ in the network does not notably change when $L_{B}<10$. In other words, even the level of most peripheral nodes typically goes up to higher than $L_{B}=10$, and only the $60$ nodes constituting the primary and secondary cores remain when $L_{B}\geq 16$. It means that the kCD roughly cuts the structure into two levels: both the primary and the secondary cores as the core and the rest as the periphery. When we apply the LED, in sharp contrast, the number of total nodes with $L\geq L_{B}$ decreases gradually, and in particular, the number of primary and secondary core nodes ($60$ nodes) gradually decreases for $L_{B}>10$, while most decrement in the number of nodes for $L_{B}<10$ is caused by the removal of peripheral nodes ($100$ nodes). Even if we set up the three-level coreness, due to the statistical fluctuation there must be differential levels of local hubs for a single realization of SBM, and the LED successfully captures them as well.
@@
We also measure $S_{cp}$ as a function of $L_{B}$, in the same manner as in Fig. 7. As shown in Fig. 11(b), in the case of the kCD, since most periphery nodes are in high levels ($>10$), $S_{cp}$ remains low across a wide range of lower $L_{B}$ values. At the maximum value of $L_{B}=19$, $S_{cp}$ reaches $0.320$ and the most densely connected $55$ nodes are classified as the core. However, in the case of the LED, $S_{cp}$ starts to increase and reaches its maximum value ($0.453$) at $L_{B}=L_{B}^{*}=16$, and decreases from that point. The number of nodes at $L_{B}^{*}$ is $20$ [Fig. 11(a)], which corresponds to the primary core nodes as expected. In other words, when using the LED, it is possible to identify the most influential nodes even within densely connected nodes. This behavior is consistent with that observed in real networks, shown in Figs. 7 and 8. As $L_{B}$ increases, the proportion of nodes belonging to $I_{1}$ increases among the nodes with $L>L_{B}$, leading to a reduction in the number of connections between nodes with $L\geq L_{B}$ and those with $L<L_{B}$. This results in a lower value of $N_{cp}/N_{p}$ and the presence of the optimal boundary level. Compared to the kCD, the changes in $S_{cp}$ values are smoother in the LED. In the real-world networks considered, many communities exhibit star or densely connected structures. In those cases, the kCD often assigns a single level to all nodes, effectively grouping them at the same level. In contrast, the LED differentiates nodes by assigning various levels on them, which naturally leads to detecting an appropriate core in that structures. Consequently, the LED has clear advantages over kCD in identifying a proper cores in these specific structures.